### v1

# Exploring Semi-Supervised Learning for Online Mapping

Abstract
Online mapping is important for scaling autonomous driving beyond well-defined areas. Training a model to produce a local map, including lane markers, road edges, and pedestrian crossings using only onboard sensory information, traditionally requires extensive labelled data, which is difficult and costly to obtain. This paper draws inspiration from semi-supervised learning techniques in other domains, demonstrating their applicability to online mapping. Additionally, we propose a simple yet effective method to exploit inherent attributes of online mapping to further enhance performance by fusing the teacher’s pseudo-labels from multiple samples. The performance gap to using all labels is reduced from 29.6 to 3.4 mIoU on Argoverse, and from 12 to 3.4 mIoU on NuScenes utilising only 10% of the labelled data. We also demonstrate strong performance in extrapolating to new cities outside those in the training data. Specifically, for challenging nuScenes, adapting from Boston to Singapore, performance increases by 6.6 mIoU when unlabelled data from Singapore is included in training.
Figure 1: Self-training with Teacher Student set-up for Online Mapping.
I Introduction
An autonomous vehicle needs to understand where it is safe to drive. This includes knowing where other road participants are and where the lane delimiters and road edges are located around the vehicle. Relying solely on pre-made offline High-Definition (HD) maps for road information comes with difficulties, such as accurate localisation within the map, and the costs associated with building and keeping the maps up to date. The latter is significant for scaling Autonomous Driving (AD) beyond limited, well-defined areas. Online mapping is an alternative and complementary approach, where onboard sensory data is used to produce a local up-to-date map depicting surrounding lane markers, road edges, and pedestrian crossings in real-time [24, 25].
Recent studies have shown that end-to-end learning-based online mapping algorithms struggle to generalise to regions not present in their training set, suggesting that the available data is insufficient for learning this complex task [27, 29]. While collecting large amounts of AD data is relatively straightforward, the bottleneck lies in the effort needed to label the data. Creating the necessary labels, e.g. 3D polylines for each lane divider, is generally time-consuming and costly. Therefore, we are interested in approaches that utilize a small amount of labelled data and a vast amount of unlabelled data to increase the generalisability in the online mapping task.
Other domains, such as image classification [39] and semantic segmentation [13], have addressed this problem setting through Semi-Supervised Learning (SSL) and shown remarkable capabilities. Recently the interest in SSL for object detection has also gained traction and shown promising results [32, 47]. Some SSL techniques have also been successfully adapted to the closely related domain of Bird’s Eye View (BEV) semantic segmentation. For instance, utilising a teacher-student setup [39] has demonstrated effectiveness in monocular BEV segmentation, applicable to both dynamic classes such as vehicles and pedestrians, and static classes including drivable areas, and walkways [55, 14]. However, new challenges arise when multiple cameras are used as input, as creating 3D-consistent augmentations in this context proves difficult. For instance, the conjoint-rotation technique introduced by Zhu et al. [55] is incompatible with our problem setup. Concurrent work has shown that randomly disabling entire camera feeds and applying dropout on the BEV features as strong augmentation for the student are effective when surround-view images are used as input [21].
In this work, we aim to identify which key SSL approaches excel in the online mapping domain and explore how to best leverage the unique characteristics of online mapping. For instance, the static nature of objects in online mapping allows for leveraging temporal consistency to enhance performance, which is not possible in the more general problem formulation of BEV semantic segmentation. In this spirit, we propose and investigate a label-efficient SSL method, illustrated in Fig. 1, that uses sequential data to enable the teacher to fuse information across multiple samples and thus produce more reliable pseudo-labels. This approach effectively boosts performance in online mapping leveraging less expensive unlabelled data.
This paper makes four key contributions: First, we conduct a comprehensive analysis of prevalent SSL techniques for online mapping, demonstrating which existing methods can significantly enhance performance. Second, we propose using the static-world assumption to improve pseudo-label accuracy by enabling the teacher model to integrate information across multiple samples. Third, we combine SSL components into a robust method that maintains high performance levels across varying volumes of labelled data. Lastly, we demonstrate this method’s additional benefit of strong performance in domain adaptation when trained in one group of cities and evaluated in another.
II Related Work
This work explores which semi-supervised learning, and domain adaptation, approaches excel when applied to the online mapping setting. Thus, we describe these fields in the following section.
II-A Semi-Supervised Learning
In SSL, we combine costly labelled data with inexpensive unlabelled data to train a model. There are two main approaches to this problem, single-stage training where automatic ground truth labels are generated during training, i.e. using predictions from a single model or an ensemble of models as pseudo-labels [37, 51, 23], or a two-stage process starting with self-supervised pre-training followed by finetuning on the labelled data [8, 19, 5]. This paper concentrates on the single-stage joint training approach, exploring critical design decisions and their impact on the performance of semi-supervised online mapping.
SSL has a rich history in image classification, with the teacher-student framework [39] being particularly influential. This framework involves a student network supervised by predictions from a teacher network. The teacher network shares the same architecture as the student but updates its weights as an exponential moving average (EMA) of the student’s weights. ReMixMatch [1] further introduced the idea of feeding the teacher model weakly augmented inputs, while the student model receives strongly augmented inputs. Many works have highlighted the importance of data augmentation in SSL [54, 13, 41] and these augmentations may differ from those suitable for fully supervised settings [54]. Common augmentation strategies for SSL include RandAugment [10], CutOut [11], CutMix [49], and feature dropout [48].
FixMatch [37] further popularized thresholding, a technique where the student is supervised only with high-confidence pseudo-labels generated by the teacher. Numerous follow-up studies continue to demonstrate the strengths of this method [48, 44]. Another line of work in SSL involves adding self-supervised auxiliary tasks to improve performance, such as predicting the parameters of stochastic image transformations applied to unlabelled images [43, 50] or adding similarity loss of the features between the teacher and student models [9, 42].
Research in SSL has also been extended to BEV segmentation, considering a variety of classes such as car, truck, road, walkway, and vegetation [14, 55, 15, 16]. The teacher-student approach has, for instance, been adapted to monocular BEV segmentation, utilizing MSE losses for probability and feature similarity [55, 14]. Unlike in Perspective View (PV) semantic segmentation, where a variety of augmentations are permissible, the augmentations applied in BEV segmentation must maintain 3D consistency. E.g., methods such as randomly cropping and resizing the PV input image for BEV segmentation can break geometric consistency in the BEV plane. One augmentation with geometric consistency is the conjoint rotation introduced in [55] where both images and BEV pseudo-labels are rotated around the upward axis. However, this technique is limited to monocular camera setups, as any augmentation affecting the 3D geometry must maintain 3D consistency across all cameras.
Our setup closely resembles the concurrent work of PCT [21], which proposes SSL on multi-view images using a teacher-student framework inspired by [48]. They introduce a camera dropout augmentation technique, where one or more camera images are omitted for the student model, and leverage BEV feature dropout. Additionally, they use pre-trained models to generate PV semantic segmentation and depth pseudo-labels. However, in this paper, we focus on core SSL methods without incorporating additional pre-trained networks for pseudo-labeling.
II-B Domain Adaptation
Unsupervised domain adaptation (UDA) is a field closely related to SSL, where the unlabelled data is drawn from a target domain that is slightly shifted from the source domain (labelled data). The goal of UDA is to achieve strong performance in the target domain and is of great importance to AD as we want to generalise to new geographical regions. UDA incorporates many techniques commonly used in SSL, with self-training being the most prominent [56]. There are many works on UDA for semantic segmentation with AD applications in mind. The dominant approach in these works involves augmenting data by mixing crops of labelled and unlabelled data [40, 7]. However, such augmentations are not well-suited for our setting, as online mapping requires consistency across multiple camera views.
II-C Online Mapping
Recently, significant progress has been made in the field of online mapping, leading to the development of two primary approaches: segmentation-based methods [46, 12, 24, 34] and vector-based methods [31, 25, 26, 6] where the key difference lies in the output representation. Segmentation-based methods rasterize a region from a Bird’s Eye View and classify each cell, typically using a CNN [35, 12, 24]. The vector-based methods predict vectors with connected points, representing a polyline for each object, typically using a DETR-like [4] transformer decoder [25, 26, 53].
Both approaches encode input images using a neural network and transform the features into BEV using a core technique known as lifting. For instance, LSS [36] learns a categorical depth distribution for each pixel and use it to weigh how much the corresponding PV feature should influence the corresponding BEV-cell.
We are mainly interested in the effects of SSL approaches and will use a segmentation-based method as the formulation is a better match for existing SSL and UDA methods.
III SSL for Online Mapping
Our work focuses on leveraging SSL to maximize the effectiveness of unlabelled data in online mapping. In this context, we assume that we have a small set of labelled data ${\mathbf{X}_{L}=\{\mathbf{x_{i}},\mathbf{y_{i}}\}_{i=1}^{n_{l}}}$, for which ground truth map $\mathbf{y}_{i}$ is available, and a large set of unlabelled images $\mathbf{X}_{U}=\{\mathbf{x_{i}}\}_{i=1}^{n_{u}}$ without ground truth annotations. The latter is assumed to be collected in areas separate from the labelled set, but using the same cameras. Further, the relative pose of the vehicle in the sequences is assumed to be known.
Using the labelled data, the student is trained in an ordinary supervised manner by minimizing a supervised loss ${\ell_{\text{sup}}(\mathbf{x_{i}},\mathbf{y_{i}})}$, e.g., cross-entropy, DICE [38], or Focal Loss [28]. Effectively using the unlabelled data is more challenging as we do not have access to a ground truth map for this area, and several design choices must be made. To understand how to apply SSL methods in the context of online mapping, we explore the impact of these choices using a well-established online mapping architecture. Figure 1 shows an overview of the SSL training scheme, and in this section, we provide the details of the online mapping architecture, the prevalent SSL techniques we employed, and our proposed simple multi-step teacher fusion approaches.
III-A Online mapping architecture
As we are mainly interested in the effects of SSL approaches, we choose to use a typical map segmentation architecture [24, 52, 34] such as Fig. 1 illustrates. A set of multi-view images, $\mathbf{x}$, is individually encoded by a neural network ${\mathbf{z}_{\text{pv}}=f^{\text{enc}}_{\theta}(\mathbf{x})}$ and lifted to BEV features as ${\mathbf{z}_{\text{bev}}=f^{\text{lift}}_{\theta}(\mathbf{\mathbf{z}_{\text{pv}}})}$. Then, another neural network decodes the BEV features ${\tilde{\mathbf{z}}_{\text{bev}}=f^{\text{dec}}_{\theta}(\mathbf{z}_{\text{bev}})}$. The final class probabilities are predicted through a classification head ${\mathbf{p}_{\text{cls}}=f^{\text{cls}}_{\theta}(\tilde{\mathbf{z}}_{\text{bev}})}$, where ${\mathbf{p}_{\text{cls}}}$ comprise the independent probability of each map grid cell containing a lane delimiter, road edge, or pedestrian crossing. As such, the online map is treated as a multi-label semantic segmentation problem. The complete architecture can be described as ${\mathbf{p}_{\text{cls}}=F_{\theta}(\mathbf{x})}$.
III-B Semi-supervised training scheme
The primary training methodology adopted in this work is based on the Mean Teacher framework [39]. As illustrated in Fig. 1, we use two identical network architectures: a teacher-version, parameterized by ${\theta_{\text{T}}}$, and a student-version, described by $\theta_{\text{S}}$. In this setup, only the student is trained directly from data, $\mathbf{X}_{L}$ and $\mathbf{X}_{U}$. The teacher weights are updated after each training step using an exponentially moving average (EMA) of the student, ${\theta_{T}\leftarrow\alpha\theta_{T}+(1-\alpha)\theta_{S}}$, where $\alpha$ is the keep rate.
We investigate two common forms of SSL schemes, pseudo-labelling and consistency regularization through feature similarity. The first follows the approach in [1, 37, 39, 56], where the student is trained using pseudo-labels (pseudo-maps for our setting) predicted by the teacher. The teacher network receives a weakly augmented version of the input images $\mathbf{x}$ (and possibly a set of additional surrounding data samples, Sec. III-F), for which it predicts a pseudo-map ${\hat{\mathbf{y}}=\varphi(F_{\theta_{T}}(\mathbf{x}))}$. Here, $\varphi(\cdot)$ is used to decide/design suitable pseudo-labels based on the predicted class probabilities, e.g., thresholding and/or sharpening (see Sec. III-D). The student receives a strongly augmented version of the same sample $\tilde{\mathbf{x}}$ (see Sec. III-C) and is trained to make consistent class predictions with the teacher model through a pseudo-label loss ${\ell_{\text{cls}}(F_{\theta_{S}}(\tilde{\mathbf{x}}),\varphi(F_{\theta_{T}}(\mathbf{x})))}$.
The feature similarity supervision on the other hand enforces consistency between intermediate features for the teacher and student models. More formally, we train the student network to minimize a feature similarity loss $\ell_{\text{feat}}(\mathbf{z}_{\theta_{T}}(\mathbf{x}),\mathbf{z}_{\theta_{S}}(\tilde{\mathbf{x}}))$, where $\mathbf{z}_{\theta_{T}}$ and $\mathbf{z}_{\theta_{S}}$ are intermediate features on the same level, e.g. $\mathbf{z}_{\text{bev}}$ or $\tilde{\mathbf{z}}_{\text{bev}}$, using teacher and student models, respectively. In contrast to the pseudo-label training, feature similarity is insensitive to errors in teacher predictions as we avoid any label decisions. We only assume that the representation of the same scene should be consistent in both models. The details of this loss are discussed in Sec. III-E.
In sum, the total training loss for the student is
$\begin{split}\ell=\sum_{\mathbf{x}_{i}\in\mathbf{X}_{L}}\ell_{\text{sup}}(\mathbf{x}_{i},\mathbf{y}_{i})&+\sum_{\mathbf{x}_{i}\in\mathbf{X}_{U}}\omega_{\text{cls}}\ell_{\text{cls}}(F_{\theta_{S}}(\tilde{\mathbf{x}}_{i}),\varphi(F_{\theta_{T}}(\mathbf{x}_{i}))\\ &+\omega_{\text{feat}}\ell_{\text{feat}}(\mathbf{z}_{\theta_{S}}(\tilde{\mathbf{x}}_{i}),\mathbf{z}_{\theta_{T}}(\mathbf{x}_{i}))),\end{split}$
where $\omega_{\text{cls}}$ and $\omega_{\text{feat}}$ are hyperparameters controlling the relative importance of the respective loss.
III-C Augmentations
The strong augmentation techniques used for the input fed to the student are crucial [54]. Thus, we investigate various data augmentation techniques that do not change the geometry of the scene, as discussed in Sec. II-A. The augmentation techniques investigated in this work include photometric augmentation (random adjustments to brightness, contrast, color conversion, saturation, hue, and channel swaps), camera drop augmentation [21], CutOut [11], and BEV feature dropout inspired by [48].
III-D Sharpening and thresholding
In SSL, predictions from the teacher are often refined through thresholding and sharpening before being used as pseudo-labels for the student. Thresholding involves keeping only predictions above some confidence level, whereas sharpening scales the predicted distribution to increase the confidence of the pseudo-label. Although there are many variations of thresholding and sharpening, we focus on a few key approaches. For thresholding, we follow FixMatch [37], i.e., keeping only predictions exceeding a fixed confidence threshold. For sharpening, we explore two techniques: (1) hard pseudo-labeling where the teacher predictions are reduced to its argmax [37, 22], and (2) temperature sharpening, where the teacher’s predicted logits are rescaled through division by a temperature parameter (denoted $T$ in [17, 2]) before applying sigmoid to get a soft pseudo-label.
III-E Feature Similarity
In addition to pseudo-labelling, we can also enforce feature similarity at the BEV feature level, i.e. adding the term $\omega_{\text{feat}}\ell_{\text{feat}}(\mathbf{z}_{\theta_{T}}(\mathbf{x}),\mathbf{z}_{\theta_{S}}(\tilde{\mathbf{x}}))$ to the total loss. This approach requires the student to maintain consistent BEV features with the teacher, which have been shown to improve performance [42]. While DoubleMatch used cosine similarity as the loss function, previous work on BEV segmentation employed Mean Squared Error (MSE) [55, 14]. We are therefore interested in exploring the impact of the chosen loss function on performance. Another design choice involves determining the optimal stage to harmonize the features between the teacher and the student. We identify two potential levels of interest: immediately following the lifting method, $\mathbf{z}_{\text{bev}}$, and just before decoding the class, $\tilde{\mathbf{z}}_{\text{bev}}$. Incorporating the feature similarity loss can potentially destabilize the training process as overemphasizing this loss may lead to a trivial solution.
III-F Teacher Fusion
Unlike BEV segmentation, which includes dynamic classes, all classes in online mapping are stationary. Since sensor data is collected in consecutive sequences with known ego vehicle movement, we can improve the teacher’s pseudo-labels by fusing predictions from multiple nearby samples. As Fig. 1 illustrates, predictions from past and future samples can be inferred and transformed into the current frame. By selecting only the most confident teacher predictions from any of the samples we allow the teacher to have a broader view of the nearby map and make more accurate pseudo-labels. This is especially powerful as predictions far from the ego vehicle tend to be less accurate [12]. As an alternative, we can fuse the BEV features across multiple samples by taking the mean of overlapping processed BEV features $\tilde{\mathbf{z}}_{\text{bev}}$.
To maximize overlap between additional samples and the area of the main sample, and to ensure that these additional samples provide new information, we select them from positions at randomly sampled distances relative to the current main frame. This approach mitigates issues associated with time-based sample selection, such as the ego vehicle being stationary—which offers no additional information—or moving at very high speeds, resulting in little overlap.
IV Experiments
To maximize the effectiveness of adding unlabelled data in the online mapping setting, we optimize parameters and design choices for each of the discussed methods and assess the relative impact of different components through ablation studies. Additionally, we examine the usefulness of our method by training on one set of cities and evaluating on another, assessing its capability for domain adaptation.
IV-A Experimental Setup
We employ a ResNet50 [20] as image encoder $f^{\text{enc}}_{\theta}$, LSS [36] for lifting features to BEV $f^{\text{lift}}_{\theta}$, and a lightweight convolutional bottleneck decoder inspired by [18] as $f^{\text{dec}}_{\theta}$. We perform multi-label prediction for pedestrian crossings, lane dividers, and road boundaries and use Focal Loss [28] as the supervised loss $\ell_{\text{sup}}$ on the labelled dataset and $\ell_{\text{cls}}$ for the unlabelled data. The region of interest follows [30] spanning ±$\pm 45$m longitudinally and ±$\pm 15$m laterally. Each cell has a side length of $0.3$m. For brevity, we report the mean Intersection over Union (mIoU) across these three classes. To account for the stochasticity in the results from individual runs, we either report the mean and standard deviation of the mIoU for three runs with the same hyperparameters or multiple nearby values of the hyperparameter. Empirically, we note that our setup is robust to a range of unlabelled weights $\omega_{\text{cls}}$ spanning $0.25$ to $5.0$. We choose to follow [14] and use $\omega_{\text{cls}}=1$. We further follow [21] and linearly ramp-up $\omega_{\text{cls}}$, and when applicable $\omega_{\text{feat}}$, over the first third of the training. Unless stated otherwise, no thresholding or sharpening is used in the experiments. The models, unless stated otherwise, are trained and evaluated using the Argoverse 2 (AV2) dataset [45] with the geographically disjoint splits proposed by [27] using surround-view camera images as input and use the provided HD maps as labels. We generally experiment using two utilisation rates of labelled data, namely $2.5\%$ and $10\%$, corresponding to $18$ and $70$ labelled sequences respectively.
IV-B Augmentations
TABLE I: Performance (mIoU) of various augmentations.
Augmentation Label utilisation
Photom. CamDrop CutOut BEVDrop $2.5\%$ $10\%$
$12.1^{\pm 0.6}$ $21.3^{\pm 0.2}$
$13.3^{\pm 0.6}$ $32.0^{\pm 0.1}$
$21.9^{\pm 1.0}$ $31.9^{\pm 0.1}$
$24.4^{\pm 1.2}$ $33.9^{\pm 0.2}$
$25.5^{\pm 0.7}$ $34.3^{\pm 0.1}$
Empirically we observe that using a high probability of applying the augmentation yields better performance and choose always to apply the strong augmentations to the student. For the CutOut augmentation, we find that removing $25\%$ of the image yields good results and is thus used for all experiments. For BEV feature dropout (BEVDrop) we follow [21, 48] and drop $50\%$ of the features. The results in Tab. I demonstrate that increasing the difficulty of the task for the student model is crucial also for online mapping. With $10\%$ label utilisation, Photometric augmentation alone is sufficient for significant model improvement. However, with $2.5\%$ label utilisation, additional augmentations are required. For both label utilisation ratios, applying Photometric, CutOut, and BEVDrop yields the best performance, and we adopt this configuration as the Baseline in further experiments. Specifically, we observe a significant performance improvement, with mIoU increasing from $12.1$ to $25.5$ for the $2.5\%$ label utilisation and $21.3$ to $34.3$ for the $10\%$ label utilization. Furthermore, contrary to the findings in [21], we find that removing parts of the images, i.e. CutOut, yields better performance than dropping an entire camera and masking its field of view in the loss, as done with CamDrop.
IV-C Sharpening and Thresholding
Fig. 2 illustrates the effects of sharpening and thresholding. As shown on the left, supervising with only the most confident pseudo-labels does not yield any additional performance boost. Using a low threshold can marginally improve performance, and we opt to use $0.6$ as the threshold in further experiments. Tab. II displays the results over multiple runs using hard and soft labels. We note that there is no significant change in performance using sharpening, which differs from other domains where sharpening has shown to be very effective, usually with high thresholds [37, 48, 44]. While applying sharpening with a temperature parameter can improve performance, right column in Fig. 2, the gains are marginal and introduce an additional hyperparameter.
Figure 2: Thresholding and sharpening utilising $2.5\%$ and $10\%$ of the labels. Baseline refers to using only strong augmentations.
TABLE II: mIoU for sharpening methods using threshold $0.6$.
Technique Label utilisation
Thr. Hard $2.5\%$ $10\%$
$25.5^{\pm 0.7}$ $34.3^{\pm 0.1}$
$25.1^{\pm 1.3}$ $33.9^{\pm 0.1}$
$25.4^{\pm 0.8}$ $34.6^{\pm 0.1}$
IV-D Feature Consistency
We compare the performance using Mean Squared Error (MSE) and Cosine Similarity (COS) losses with $\omega_{\text{cls}}=1$ and varying feature loss weight $\omega_{\text{feat}}$. In Fig. 3 the results are displayed using features from before (Pre) or after (Post) the BEV decoder $f^{\text{dec}}_{\theta}$ as inputs to $\ell_{\text{feat}}$. When utilizing $2.5\%$ of the labels the results are noisy, and no clear benefit from the additional supervision is observed across the setups. However, for the $10\%$ label utilisation, COS consistency provides a slight performance gain and demonstrates greater stability than the MSE loss. While prior BEV segmentation works opted to use MSE [55, 14], we believe COS to be the better choice.
Figure 3: Feature Similarity using Mean Square Error and Cosine Similarity with varying weights. Baseline refers to using only strong augmentations. Pre and Post refers to using $\mathbf{z}_{\text{bev}}$ or $\mathbf{\tilde{z}}_{\text{bev}}$ respectively.
IV-E Teacher Fusion
To enhance pseudo-labels, we fuse teacher predictions from multiple samples. We use two additional samples, as we empirically found that adding more samples does not provide further benefits. Therefore, we focus on examining the influence of the maximum distance to these additional samples. The results in Fig. 4 show that even our simple fusion of predicted probabilities generally improves performance, both with and without thresholding. The fusion of BEV features, on the other hand, shows no significant benefit. For $2.5\%$ label utilisation we observe that increasing the range to above $20$m is required to improve performance, whereas for $10\%$ the maximum range is not as important. In further experiments, we will use $30$m as it works well for both utilisation ratios. While the simple probability fusion strategy yields strong results, exploring more advanced fusion techniques is a promising avenue for future research.
Figure 4: Teacher Fusion across varying maximum ranges from the current sample. Baseline refers using only strong augmentations.
IV-F Combining Components
Building on the above results, we further explore how these components interact in SSL for online mapping. Starting with the teacher-student setup as a Core we incrementally add components and evaluate their impact on performance. Here, we add strong augmentations to the student, +Augs, with Photometric, CutOut, and BEVDrop. Teacher fusion, +Fusion, is used with a maximum range of $30$m. The Cosine feature similarity loss is incorporated under +Featsim with a weight of $w_{\text{feat}}=0.25$. Additionally, a threshold of $0.6$ is used without any temperature sharpening in +Thr, and hard pseudo-labeling is further added and denoted with +Hard.
The results in Tab. III demonstrate that combining components leads to further increased performance. Adding strong augmentations to the student yields a significant performance boost, as does further incorporating teacher fusion. Together these contribute to mIoU gains of $16.1$ and $15.1$ for $2.5\%$ and $10\%$ respectively. When combined with Fusion, feature similarity slightly improves performance for $2.5\%$ labelled data, but has no impact for $10\%$. Finally, applying thresholding yields a modest improvement for both ratios, while using the hard pseudo-labels does not contribute further. These results underscore the importance of conducting ablation studies across different labelled data ratios, as the impact of various modules can vary depending on the amount of labelled data.
TABLE III: mIoU when incrementally adding SSL components.
Label Util. Core +Augs +Fusion +Featsim +Thr. +Hard
$2.5\%$ $12.1^{\pm 0.6}$ $25.5^{\pm 0.7}$ $28.2^{\pm 0.7}$ $29.0^{\pm 0.6}$ $29.4^{\pm 0.3}$ $29.3^{\pm 1.1}$
$10\%$ $21.3^{\pm 0.2}$ $34.3^{\pm 0.1}$ $36.4^{\pm 0.2}$ $36.4^{\pm 0.5}$ $36.8^{\pm 0.2}$ $36.7^{\pm 0.3}$
IV-G Varying Label Utilisation
We are also interested in how our proposed SSL method scales with varying amounts of labelled data. Using the parameters from Sec. IV-F, our proposed method includes strong augmentations, teacher fusion, Cosine feature similarity, and thresholding. Tab. IV displays performance when using both unlabelled and labelled data against using only the labelled data. Each setup is repeated three times with different data folds. As before, we report the mean and standard deviation; however, in this case, we select the best-performing model on the validation set and evaluate it on the test set. Besides conducting experiments on AV2 [45], we also use NuScenes [3], using the same hyperparameters except for learning rate and weight decay. As expected, performance decreases as the amount of labelled data is reduced. Here, we experiment using $100\%$ of the data, corresponding to $700$ sequences, down to only $18$ sequences at $2.5\%$ utilisation. Using the unlabelled data and our proposed setup yields significant performance gains across all labelled data utilisation ratios. In the lowest setting, relying solely on the labelled sequences results in poor performance, with mIoU dropping as low as $7.5$ for AV2 and $5.6$ for nuScenes. However, incorporating unlabelled data significantly boosts performance to $28.9$ and $15.5$ mIoU for the two datasets respectively. Using only $10\%$ of the labels, the performance gap between using the unlabelled data and not is reduced from $29.6$ mIoU to $3.5$ mIoU on AV2, and from $12$ mIoU to $3.4$ mIoU on NuScenes.
TABLE IV: Test performance (mIoU) for multiple label utilisations for both Argoverse 2 (AV2) and NuScenes (NuSc).
+Unlbld Label utilisation
Data $2.5\%$ $5\%$ $10\%$ $25\%$ $50\%$ $100\%$
AV2 $7.5^{\pm 0.9}$ $9.4^{\pm 0.1}$ $10.6^{\pm 0.2}$ $23.4^{\pm 0.5}$ $33.8^{\pm 0.7}$ $40.2^{\pm 0.7}$
$\mathbf{28.9}^{\pm 1.2}$ $\mathbf{33.8}^{\pm 0.6}$ $\mathbf{36.7}^{\pm 0.9}$ $\mathbf{37.4}^{\pm 0.8}$ $\mathbf{39.6}^{\pm 0.3}$-
NuSc $5.6^{\pm 0.3}$ $7.2^{\pm 2.3}$ $16.6^{\pm 3.0}$ $24.5^{\pm 2.2}$ $27.4^{\pm 2.7}$ $28.6^{\pm 2.4}$
$\mathbf{15.5}^{\pm 1.2}$ $\mathbf{22.0}^{\pm 1.6}$ $\mathbf{25.1}^{\pm 1.9}$ $\mathbf{27.4}^{\pm 1.8}$ $27.8^{\pm 2.8}$-
IV-H City Adaptation
We are interested in determining whether the proposed method can be used to adapt the model to a new city. To investigate this, we evaluate the model’s performance using the AV2 Far Extrapolation splits proposed by [27] and the Boston to Singapore split used in [33] for NuScenes. For all runs, we use all labelled data from the source cities while gradually increasing the amount of unlabelled data from the target cities, ensuring that this data remains geographically disjoint from the test set. As reported in Tab. V, we find that, despite having no labelled data from the target city, we can significantly improve performance in that city by leveraging unlabelled data from it. In the AV2 dataset, our approach nearly matches the performance of a model trained using all target labels not included in the test set. E.g., when adapting to Pittsburgh, the performance difference between using labelled data and our method is only $0.5$ mIoU. However, it is noteworthy that the domain gap between the cities in AV2 is smaller since all cities are located in the United States. For NuScenes, the domain gap between Boston and Singapore is substantial, making it a more challenging problem. Yet our proposed method increases the mIoU from $11.1$ to $17.0$ using $100$ unlabelled sequences from Singapore.
TABLE V: Performance (mIoU) for city adaptaion for both Argoverse 2 and NuScenes when increasing amount of unlabelled data.
Folds # unlabelled target sequences Using target
00 $25$ $50$ $100$ $200$ labels
AV2 PIT-MIA $\rightarrow$ Rest $33.6$ $34.4$ $35.5$ $37.5$ $37.2$ $38.9$
MIA-Rest $\rightarrow$ PIT $36.3$ $38.7$ $40.0$ $40.2$ $40.7$ $41.2$
PIT-Rest $\rightarrow$ Mia $34.5$ $36.1$ $37.5$ $37.8$ $38.5$ $40.2$
NuSc Boston $\rightarrow$ Singapore $11.1$ $12.6$ $15.3$ $17.0$ $17.7$ $27.0$
V Conclusion
This paper draws inspiration from semi-supervised learning (SSL) techniques in other domains, demonstrating how to apply them to online mapping. For instance, we have shown that applying Photometric, CutOut, and BEV feature dropout augmentations to the student network significantly improves performance in a teacher-student setup. Further, we show that our proposed multi-step fusion for improving the teachers’ pseudo-labels yields an additional performance boost and that the Cosine similarity loss can help improve learning.
Combining the methods explored in this paper, performance can be significantly increased when also utilising unlabelled data for both Argoverse 2 and NuScenes compared to only using the labelled data. For instance, merely using $10\%$ of the labels, the performance gap between using the unlabelled data and not is reduced from $29.6$ mIoU to $3.5$ mIoU on Argoverse 2, and from $12$ mIoU to $3.4$ mIoU on NuScenes. We have also shown that our method can be used for city adaptation, where we have labels for one city but not for the other. For example, when adapting from Boston to Singapore, performance is increased by $6.6$ mIoU when unlabelled data from Singapore is included during training.
In summary, efficient use of both labelled and unlabelled data is crucial for advancing the field of online mapping. While we could not explore all existing SSL methods, this work takes a step toward understanding the strengths and weaknesses of core SSL methods in this context, aiming to inspire further research in this area. Future research could explore more rigorous teacher fusion over multiple samples or incorporate self-supervision tasks for further advancement.
Acknowledgements We want to express our sincere gratitude to Adam Tonderski, Georg Hess, William Ljungbergh, Carl Lindström, and Ji Lan for their continuous support and invaluable discussions. This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at NSC Berzelius and C3SE Alvis partially funded by the Swedish Research Council through grant agreement no. 2022-06725.

### v2

@@
 Abstract
-Online mapping is important for scaling autonomous driving beyond well-defined areas.
-Training a model to produce a local map, including lane markers, road edges, and pedestrian crossings using only onboard sensory information, traditionally requires extensive labelled data, which is difficult and costly to obtain.
-This paper draws inspiration from semi-supervised learning techniques in other domains, demonstrating their applicability to online mapping.
-Additionally, we propose a simple yet effective method to exploit inherent attributes of online mapping to further enhance performance by fusing the teacher’s pseudo-labels from multiple samples.
-The performance gap to using all labels is reduced from 29.6 to 3.4 mIoU on Argoverse, and from 12 to 3.4 mIoU on NuScenes utilising only 10% of the labelled data.
-We also demonstrate strong performance in extrapolating to new cities outside those in the training data.
-Specifically, for challenging nuScenes, adapting from Boston to Singapore, performance increases by 6.6 mIoU when unlabelled data from Singapore is included in training.
-Figure 1: Self-training with Teacher Student set-up for Online Mapping.
-I Introduction
-An autonomous vehicle needs to understand where it is safe to drive.
-This includes knowing where other road participants are and where the lane delimiters and road edges are located around the vehicle.
-Relying solely on pre-made offline High-Definition (HD) maps for road information comes with difficulties, such as accurate localisation within the map, and the costs associated with building and keeping the maps up to date.
-The latter is significant for scaling Autonomous Driving (AD) beyond limited, well-defined areas.
-Online mapping is an alternative and complementary approach, where onboard sensory data is used to produce a local up-to-date map depicting surrounding lane markers, road edges, and pedestrian crossings in real-time [24, 25].
-Recent studies have shown that end-to-end learning-based online mapping algorithms struggle to generalise to regions not present in their training set, suggesting that the available data is insufficient for learning this complex task [27, 29].
-While collecting large amounts of AD data is relatively straightforward, the bottleneck lies in the effort needed to label the data.
+The ability to generate online maps using only onboard sensory information is crucial for enabling autonomous driving beyond well-mapped areas.
+Training models for this task – predicting lane markers, road edges, and pedestrian crossings – traditionally require extensive labelled data, which is expensive and labour-intensive to obtain.
+While semi-supervised learning (SSL) has shown promise in other domains, its potential for online mapping remains largely underexplored.
+In this work, we bridge this gap by demonstrating the effectiveness of SSL methods for online mapping.
+Furthermore, we introduce a simple yet effective method leveraging the inherent properties of online mapping by fusing the teacher’s pseudo-labels from multiple samples, enhancing the reliability of self-supervised training.
+If 10% of the data has labels, our method to leverage unlabelled data achieves a 3.5x performance boost compared to only using the labelled data.
+This narrows the gap to a fully supervised model, using all labels, to just 3.5 mIoU.
+We also show strong generalization to unseen cities.
+Specifically, in Argoverse 2, when adapting to Pittsburgh, incorporating purely unlabelled target-domain data reduces the performance gap from 5 to 0.5 mIoU.
+These results highlight the potential of SSL as a powerful tool for solving the online mapping problem, significantly reducing reliance on labelled data.
+Figure 1: Semi-supervised learning framework for online mapping with temporal pseudo-label fusion.
+Our approach adapts the Teacher-Student paradigm to the online mapping domain while introducing a novel temporal fusion mechanism that aggregates pseudo-labels across multiple frames.
+We enable label-efficient learning from largely unlabelled datasets and strong generalization to unseen cities.
+1 Introduction
+Safe autonomous navigation requires accurately perceiving both dynamic traffic participants and static road infrastructure.
+Relying solely on pre-made offline High-Definition (HD) maps for the latter comes with difficulties, such as accurate localisation within the map, and the costs associated with building and keeping the maps up to date.
+The latter is significant for scaling Autonomous Driving (AD) beyond limited, well-mapped areas.
+Online mapping is an alternative and complementary approach, where onboard sensory data is used to produce a local up-to-date map of surrounding lane markers, road edges, and pedestrian crossings in real-time [25, 26].
+Recent studies have shown that learning-based online mapping algorithms struggle to generalise to regions outside their training set, suggesting available data is insufficient for learning this complex task [28, 30, 40].
+While collecting large amounts of AD data is relatively straightforward, the bottleneck lies in the effort labelling the data.
 Creating the necessary labels, e.g.
 3D polylines for each lane divider, is generally time-consuming and costly.
-Therefore, we are interested in approaches that utilize a small amount of labelled data and a vast amount of unlabelled data to increase the generalisability in the online mapping task.
-Other domains, such as image classification [39] and semantic segmentation [13], have addressed this problem setting through Semi-Supervised Learning (SSL) and shown remarkable capabilities.
-Recently the interest in SSL for object detection has also gained traction and shown promising results [32, 47].
-Some SSL techniques have also been successfully adapted to the closely related domain of Bird’s Eye View (BEV) semantic segmentation.
-For instance, utilising a teacher-student setup [39] has demonstrated effectiveness in monocular BEV segmentation, applicable to both dynamic classes such as vehicles and pedestrians, and static classes including drivable areas, and walkways [55, 14].
-However, new challenges arise when multiple cameras are used as input, as creating 3D-consistent augmentations in this context proves difficult.
-For instance, the conjoint-rotation technique introduced by Zhu et al.
-[55] is incompatible with our problem setup.
-Concurrent work has shown that randomly disabling entire camera feeds and applying dropout on the BEV features as strong augmentation for the student are effective when surround-view images are used as input [21].
-In this work, we aim to identify which key SSL approaches excel in the online mapping domain and explore how to best leverage the unique characteristics of online mapping.
-For instance, the static nature of objects in online mapping allows for leveraging temporal consistency to enhance performance, which is not possible in the more general problem formulation of BEV semantic segmentation.
-In this spirit, we propose and investigate a label-efficient SSL method, illustrated in Fig.
-1, that uses sequential data to enable the teacher to fuse information across multiple samples and thus produce more reliable pseudo-labels.
-This approach effectively boosts performance in online mapping leveraging less expensive unlabelled data.
-This paper makes four key contributions: First, we conduct a comprehensive analysis of prevalent SSL techniques for online mapping, demonstrating which existing methods can significantly enhance performance.
-Second, we propose using the static-world assumption to improve pseudo-label accuracy by enabling the teacher model to integrate information across multiple samples.
-Third, we combine SSL components into a robust method that maintains high performance levels across varying volumes of labelled data.
-Lastly, we demonstrate this method’s additional benefit of strong performance in domain adaptation when trained in one group of cities and evaluated in another.
-II Related Work
-This work explores which semi-supervised learning, and domain adaptation, approaches excel when applied to the online mapping setting. Thus, we describe these fields in the following section.
-II-A Semi-Supervised Learning
-In SSL, we combine costly labelled data with inexpensive unlabelled data to train a model.
-There are two main approaches to this problem, single-stage training where automatic ground truth labels are generated during training, i.e.
-using predictions from a single model or an ensemble of models as pseudo-labels [37, 51, 23], or a two-stage process starting with self-supervised pre-training followed by finetuning on the labelled data [8, 19, 5].
+Therefore, we are interested in approaches utilizing a small amount of labelled data and a vast amount of unlabelled data to increase the generalisability in the online mapping task.
+Other domains, such as image classification [43] and semantic segmentation [14], have addressed this problem setting through Semi-Supervised Learning (SSL) and shown strong capabilities.
+Recently, the interest in SSL for object detection has also gained traction and shown promising results [33, 51].
+Some SSL techniques have been successfully adapted to the closely related domain of Bird’s Eye View (BEV) semantic segmentation of the combination of dynamic classes such as vehicles and pedestrians, and static classes including vegetation and drivable areas.
+However, the focus has been on monocular BEV segmentation where the teacher-student setup [43] has shown effectiveness [60, 15].
+New challenges arise when multiple cameras are used as input, as creating 3D-consistent augmentations in this context proves difficult.
+For instance, the conjoint-rotation technique [60] is incompatible with our problem setup.
+Concurrent work [22] shows that randomly disabling camera feeds and applying BEV feature dropout as student augmentation techniques effectively enhance performance in a multi-camera setup.
+Despite these advances, there remains a gap in understanding how SSL components specifically interact in the context of online mapping, where the focus is exclusively on static classes rather than the mixed dynamic-static scenario addressed in prior work.
+In this work, we identify key SSL approaches that excel in online mapping, both for learning with limited labels and for adapting models to new cities.
+As online mapping focuses solely on static classes, it allows us to exploit temporal consistency across frames.
+Building on this insight, we propose a label-efficient SSL method, illustrated in Fig.
+1, that adapts the Teacher-Student paradigm to online mapping while introducing a novel temporal fusion mechanism.
+This mechanism aggregates pseudo-labels across multiple frames, improving their reliability and reducing noise.
+By leveraging large-scale unlabelled datasets, our approach significantly reduces dependence on labelled data while enabling strong generalization to unseen cities.
+To summarize, we make four key contributions: 1) a comprehensive analysis of prevalent SSL techniques for online mapping, demonstrating which existing methods can significantly enhance performance.
+2) combine SSL components into a robust method that maintains high performance levels across varying volumes of labelled data.
+3) propose that the teacher model integrates information across multiple samples to improve pseudo-label accuracy.
+4) demonstrate strong performance in domain adaptation when trained in one group of cities and evaluated in another.
+2 Related Work
+We explore semi-supervised learning and domain adaptation techniques specifically applied to online mapping. This section examines relevant prior work in these fields.
+2.1 Semi-Supervised Learning
+In SSL, costly labelled data is combined with inexpensive unlabelled data to train a model.
+There are two main approaches to this problem: single-stage training, where automatic ground truth labels are generated during training, i.e.
+using predictions from a single model or an ensemble of models as pseudo-labels [41, 55, 24], or a two-stage process starting with self-supervised pre-training followed by finetuning on the labelled data [9, 21, 5, 34].
 This paper concentrates on the single-stage joint training approach, exploring critical design decisions and their impact on the performance of semi-supervised online mapping.
@@
 The teacher network shares the same architecture as the student but updates its weights as an exponential moving average (EMA) of the student’s weights.
-ReMixMatch [1] further introduced the idea of feeding the teacher model weakly augmented inputs, while the student model receives strongly augmented inputs.
-Many works have highlighted the importance of data augmentation in SSL [54, 13, 41] and these augmentations may differ from those suitable for fully supervised settings [54].
-Common augmentation strategies for SSL include RandAugment [10], CutOut [11], CutMix [49], and feature dropout [48].
+ReMixMatch [1] further feeds the teacher weakly augmented inputs, while the student receives strongly augmented inputs.
+Data augmentation is crucial in SSL [58, 14, 45] and these may differ from those suitable for fully supervised settings [58].
+Commonly used augmentation strategies include RandAugment [11], CutOut [12], CutMix [53], and feature dropout [52].
 FixMatch [41] further popularized thresholding, a technique where the student is supervised only with high-confidence pseudo-labels generated by the teacher.
@@
 Our setup closely resembles the concurrent work of PCT [22], which proposes SSL on multi-view images using a teacher-student framework inspired by [52].
-They introduce a camera dropout augmentation technique, where one or more camera images are omitted for the student model, and leverage BEV feature dropout.
+They introduce camera dropout augmentation, where one or more camera images are omitted for the student model, and leverage BEV feature dropout.
 Additionally, they use pre-trained models to generate PV semantic segmentation and depth pseudo-labels.
-However, in this paper, we focus on core SSL methods without incorporating additional pre-trained networks for pseudo-labeling.
-II-B Domain Adaptation
-Unsupervised domain adaptation (UDA) is a field closely related to SSL, where the unlabelled data is drawn from a target domain that is slightly shifted from the source domain (labelled data).
-The goal of UDA is to achieve strong performance in the target domain and is of great importance to AD as we want to generalise to new geographical regions.
-UDA incorporates many techniques commonly used in SSL, with self-training being the most prominent [56].
-There are many works on UDA for semantic segmentation with AD applications in mind.
-The dominant approach in these works involves augmenting data by mixing crops of labelled and unlabelled data [40, 7].
-However, such augmentations are not well-suited for our setting, as online mapping requires consistency across multiple camera views.
-II-C Online Mapping
+In contrast, our work focuses on SSL methods that do not rely on additional pre-trained networks for pseudo-labeling.
+Another key distinction is that, unlike BEV segmentation, which includes dynamic object classes, online mapping is concerned solely with stationary elements.
+2.2 Domain Adaptation
+Unsupervised domain adaptation (UDA) is closely related to SSL, and throughout this paper, we will use the terms interchangeably.
+In UDA, the unlabelled data comes from a target domain that differs from the labelled source domain.
+The goal of UDA is to achieve strong performance in unseen environments, which is particularly critical for online mapping, where models must perform well in cities they have never been trained on – a problem often overlooked in prior work [28].
+Similar to SSL, many UDA techniques rely on self-training with a teacher-student setup and have also demonstrated strong results across various UDA settings [61].
+A dominant approach for perspective view semantic segmentation in the image plane for AD involves augmenting data by mixing crops of labelled and unlabelled samples [44, 7].
+However, such augmentations are poorly suited for online mapping, as they fail to maintain consistency across multiple camera views, a key requirement in our setting.
+In this work, we focus on self-training for UDA in online mapping, emphasizing its importance in adapting to unseen cities – a fundamental challenge for real-world deployment.
+2.3 Online Mapping
 Recently, significant progress has been made in the field of online mapping, leading to the development of two primary approaches: segmentation-based methods [50, 13, 25, 37] and vector-based methods [32, 26, 27, 6] where the key difference lies in the output representation.
@@
 For instance, LSS [39] learns a categorical depth distribution for each pixel and use it to weigh how much the corresponding PV feature should influence the corresponding BEV-cell.
-We are mainly interested in the effects of SSL approaches and will use a segmentation-based method as the formulation is a better match for existing SSL and UDA methods.
-III SSL for Online Mapping
+Previous works have highlighted a performance drop as the domain gap between training and evaluation data increases.
+Particularly in the context of geographical distribution shifts [28, 40], and driving scenario shifts [30].
+Our focus is on the effects of SSL approaches, and for this work, we choose a segmentation-based method, as its formulation aligns more naturally with existing SSL and UDA techniques.
+3 SSL for Online Mapping
 Our work focuses on leveraging SSL to maximize the effectiveness of unlabelled data in online mapping.
 In this context, we assume that we have a small set of labelled data ${\mathbf{X}_{L}=\{\mathbf{x_{i}},\mathbf{y_{i}}\}_{i=1}^{n_{l}}}$, for which ground truth map $\mathbf{y}_{i}$ is available, and a large set of unlabelled images $\mathbf{X}_{U}=\{\mathbf{x_{i}}\}_{i=1}^{n_{u}}$ without ground truth annotations.
-The latter is assumed to be collected in areas separate from the labelled set, but using the same cameras.
+The latter is assumed to be collected in areas separate from the labelled set, but using the same camera setup.
 Further, the relative pose of the vehicle in the sequences is assumed to be known.
@@
 Effectively using the unlabelled data is more challenging as we do not have access to a ground truth map for this area, and several design choices must be made.
-To understand how to apply SSL methods in the context of online mapping, we explore the impact of these choices using a well-established online mapping architecture.
-Figure 1 shows an overview of the SSL training scheme, and in this section, we provide the details of the online mapping architecture, the prevalent SSL techniques we employed, and our proposed simple multi-step teacher fusion approaches.
-III-A Online mapping architecture
+To understand how to apply SSL methods in the context of online mapping, we explore the impact of these choices using a well-established segmentation-based online mapping architecture.
+Figure 1 shows an overview of the SSL training scheme, and in this section, we provide the details of the online mapping architecture, the prevalent SSL techniques we employed, and our proposed multi-step teacher fusion approaches.
+3.1 Online mapping architecture
 As we are mainly interested in the effects of SSL approaches, we choose to use a typical map segmentation architecture [25, 56, 37] such as Fig.
 1 illustrates.
-A set of multi-view images, $\mathbf{x}$, is individually encoded by a neural network ${\mathbf{z}_{\text{pv}}=f^{\text{enc}}_{\theta}(\mathbf{x})}$ and lifted to BEV features as ${\mathbf{z}_{\text{bev}}=f^{\text{lift}}_{\theta}(\mathbf{\mathbf{z}_{\text{pv}}})}$.
+The multi-view images, $\mathbf{x}$, are encoded by a neural network ${\mathbf{z}_{\text{pv}}=f^{\text{enc}}_{\theta}(\mathbf{x})}$ and lifted to BEV features as ${\mathbf{z}_{\text{bev}}=f^{\text{lift}}_{\theta}(\mathbf{\mathbf{z}_{\text{pv}}})}$.
 Then, another neural network decodes the BEV features ${\tilde{\mathbf{z}}_{\text{bev}}=f^{\text{dec}}_{\theta}(\mathbf{z}_{\text{bev}})}$.
-The final class probabilities are predicted through a classification head ${\mathbf{p}_{\text{cls}}=f^{\text{cls}}_{\theta}(\tilde{\mathbf{z}}_{\text{bev}})}$, where ${\mathbf{p}_{\text{cls}}}$ comprise the independent probability of each map grid cell containing a lane delimiter, road edge, or pedestrian crossing.
+Final class probabilities are predicted through a classification head ${\mathbf{p}_{\text{cls}}=f^{\text{cls}}_{\theta}(\tilde{\mathbf{z}}_{\text{bev}})}$, where ${\mathbf{p}_{\text{cls}}}$ comprise the independent probability of each map grid cell containing a lane delimiter, road edge, or pedestrian crossing.
 As such, the online map is treated as a multi-label semantic segmentation problem.
 The complete architecture can be described as ${\mathbf{p}_{\text{cls}}=F_{\theta}(\mathbf{x})}$.
-III-B Semi-supervised training scheme
+3.2 Semi-supervised training scheme
 The primary training methodology adopted in this work is based on the Mean Teacher framework [43].
@@
 The teacher network receives a weakly augmented version of the input images $\mathbf{x}$ (and possibly a set of additional surrounding data samples, Sec.
-III-F), for which it predicts a pseudo-map ${\hat{\mathbf{y}}=\varphi(F_{\theta_{T}}(\mathbf{x}))}$.
+3.6), for which it predicts a pseudo-map ${\hat{\mathbf{y}}=\varphi(F_{\theta_{T}}(\mathbf{x}))}$.
 Here, $\varphi(\cdot)$ is used to decide/design suitable pseudo-labels based on the predicted class probabilities, e.g., thresholding and/or sharpening (see Sec.
-III-D).
+3.4).
 The student receives a strongly augmented version of the same sample $\tilde{\mathbf{x}}$ (see Sec.
-III-C) and is trained to make consistent class predictions with the teacher model through a pseudo-label loss ${\ell_{\text{cls}}(F_{\theta_{S}}(\tilde{\mathbf{x}}),\varphi(F_{\theta_{T}}(\mathbf{x})))}$.
-The feature similarity supervision on the other hand enforces consistency between intermediate features for the teacher and student models.
-More formally, we train the student network to minimize a feature similarity loss $\ell_{\text{feat}}(\mathbf{z}_{\theta_{T}}(\mathbf{x}),\mathbf{z}_{\theta_{S}}(\tilde{\mathbf{x}}))$, where $\mathbf{z}_{\theta_{T}}$ and $\mathbf{z}_{\theta_{S}}$ are intermediate features on the same level, e.g.
+3.3) and is trained to make consistent class predictions with the teacher model through a pseudo-label loss ${\ell_{\text{cls}}(F_{\theta_{S}}(\tilde{\mathbf{x}}),\varphi(F_{\theta_{T}}(\mathbf{x})))}$ similar to $\ell_{\text{sup}}$.
+The feature similarity supervision, on the other hand, enforces consistency between intermediate features for the teacher and student models.
+More formally, we train the student network to minimize a feature similarity loss, using e.g.
+Mean Squared Error or Cosine Similarity, $\ell_{\text{feat}}(\mathbf{z}_{\theta_{T}}(\mathbf{x}),\mathbf{z}_{\theta_{S}}(\tilde{\mathbf{x}}))$.
+Here, $\mathbf{z}_{\theta_{T}}$ and $\mathbf{z}_{\theta_{S}}$ are intermediate features on the same level, e.g.
 $\mathbf{z}_{\text{bev}}$ or $\tilde{\mathbf{z}}_{\text{bev}}$, using teacher and student models, respectively.
 In contrast to the pseudo-label training, feature similarity is insensitive to errors in teacher predictions as we avoid any label decisions.
-We only assume that the representation of the same scene should be consistent in both models.
-The details of this loss are discussed in Sec.
-III-E.
+Instead, we enforce the representation of the same scene to be consistent in both models and is further discussed in Sec.
+3.5.
 In sum, the total training loss for the student is
@@
 where $\omega_{\text{cls}}$ and $\omega_{\text{feat}}$ are hyperparameters controlling the relative importance of the respective loss.
-III-C Augmentations
-The strong augmentation techniques used for the input fed to the student are crucial [54].
-Thus, we investigate various data augmentation techniques that do not change the geometry of the scene, as discussed in Sec.
-II-A.
-The augmentation techniques investigated in this work include photometric augmentation (random adjustments to brightness, contrast, color conversion, saturation, hue, and channel swaps), camera drop augmentation [21], CutOut [11], and BEV feature dropout inspired by [48].
-III-D Sharpening and thresholding
-In SSL, predictions from the teacher are often refined through thresholding and sharpening before being used as pseudo-labels for the student.
-Thresholding involves keeping only predictions above some confidence level, whereas sharpening scales the predicted distribution to increase the confidence of the pseudo-label.
-Although there are many variations of thresholding and sharpening, we focus on a few key approaches.
-For thresholding, we follow FixMatch [37], i.e., keeping only predictions exceeding a fixed confidence threshold.
-For sharpening, we explore two techniques: (1) hard pseudo-labeling where the teacher predictions are reduced to its argmax [37, 22], and (2) temperature sharpening, where the teacher’s predicted logits are rescaled through division by a temperature parameter (denoted $T$ in [17, 2]) before applying sigmoid to get a soft pseudo-label.
-III-E Feature Similarity
-In addition to pseudo-labelling, we can also enforce feature similarity at the BEV feature level, i.e.
-adding the term $\omega_{\text{feat}}\ell_{\text{feat}}(\mathbf{z}_{\theta_{T}}(\mathbf{x}),\mathbf{z}_{\theta_{S}}(\tilde{\mathbf{x}}))$ to the total loss.
-This approach requires the student to maintain consistent BEV features with the teacher, which have been shown to improve performance [42].
-While DoubleMatch used cosine similarity as the loss function, previous work on BEV segmentation employed Mean Squared Error (MSE) [55, 14].
-We are therefore interested in exploring the impact of the chosen loss function on performance.
-Another design choice involves determining the optimal stage to harmonize the features between the teacher and the student.
-We identify two potential levels of interest: immediately following the lifting method, $\mathbf{z}_{\text{bev}}$, and just before decoding the class, $\tilde{\mathbf{z}}_{\text{bev}}$.
-Incorporating the feature similarity loss can potentially destabilize the training process as overemphasizing this loss may lead to a trivial solution.
-III-F Teacher Fusion
-Unlike BEV segmentation, which includes dynamic classes, all classes in online mapping are stationary.
-Since sensor data is collected in consecutive sequences with known ego vehicle movement, we can improve the teacher’s pseudo-labels by fusing predictions from multiple nearby samples.
-As Fig.
-1 illustrates, predictions from past and future samples can be inferred and transformed into the current frame.
-By selecting only the most confident teacher predictions from any of the samples we allow the teacher to have a broader view of the nearby map and make more accurate pseudo-labels.
-This is especially powerful as predictions far from the ego vehicle tend to be less accurate [12].
-As an alternative, we can fuse the BEV features across multiple samples by taking the mean of overlapping processed BEV features $\tilde{\mathbf{z}}_{\text{bev}}$.
-To maximize overlap between additional samples and the area of the main sample, and to ensure that these additional samples provide new information, we select them from positions at randomly sampled distances relative to the current main frame.
-This approach mitigates issues associated with time-based sample selection, such as the ego vehicle being stationary—which offers no additional information—or moving at very high speeds, resulting in little overlap.
-IV Experiments
-To maximize the effectiveness of adding unlabelled data in the online mapping setting, we optimize parameters and design choices for each of the discussed methods and assess the relative impact of different components through ablation studies.
-Additionally, we examine the usefulness of our method by training on one set of cities and evaluating on another, assessing its capability for domain adaptation.
-IV-A Experimental Setup
+3.3 Augmentations
+Strong augmentation techniques applied to the student’s input are essential for effective learning [58].
+Therefore, we explore various data augmentation strategies that preserve the scene’s geometry, as discussed in Sec.
+2.1.
+Specifically, we employ photometric augmentations, CamDrop [22], CutOut [12], and BEVDrop [52].
+Photometric augmentations introduce random variations in appearance by adjusting brightness, contrast, saturation, hue, and color conversion, as well as performing random channel swaps.
+CamDrop augmentation simulates missing camera views by completely removing the input from one or more cameras.
+The dropped camera’s field of view is also masked from the loss.
+CutOut randomly masks out rectangular regions in the input images while BEVDrop operates on the BEV feature representation randomly setting certain grid cells to zero.
+3.4 Sharpening and thresholding
+Predictions from the teacher are often refined through thresholding and sharpening before being used as pseudo-labels for the student.
+Thresholding keeps only predictions above some confidence level, whereas sharpening scales the predicted distribution to alter the confidence of the pseudo-label.
+For thresholding, we adapt FixMatch [41], i.e., only predictions exceeding a fixed confidence threshold contribute to the loss.
+For sharpening, we explore two techniques: (1) hard pseudo-labeling where teacher predictions are reduced to its argmax [41, 23], and (2) temperature sharpening, where the predicted logits are rescaled through division by a temperature parameter (denoted $T$ in [18, 2]) before applying sigmoid to get a soft pseudo-label.
+More formally, splitting up the classification head, $\mathbf{p}_{\text{cls}}=f^{\text{cls}}_{\theta}(\mathbf{\tilde{z}}_{\text{bev}})=\sigma(f^{\text{logits}}_{\theta}(\tilde{\mathbf{z}}_{\text{bev})})$, we have the intermediate logits $\mathbf{\hat{z}}$.
+Sharpening using the hyperparameter $T$ is then defined as $\mathbf{\hat{z}}_{sharp}=\mathbf{\hat{z}}/T$.
+For $10<T<$ 1 the certainty is artificially increased, i.e.
+$\mathbf{\hat{z}}_{sharp}>\mathbf{\hat{z}}$.
+Conversely, $T>1$ decreases the pseudo-label certainty.
+3.5 Feature Similarity
+To further enhance consistency between the teacher and student networks, we impose feature similarity constraints at the BEV feature level.
+A key consideration in this approach is selecting both the loss function and the feature level at which to apply the constraint.
+We explore two candidate feature levels for enforcing similarity: Early BEV representation ($\mathbf{z}_{\text{bev}}$) – immediately after the lifting process, capturing the initial transformation of multi-view image features into the BEV space.
+Late BEV representation ($\tilde{\mathbf{z}}_{\text{bev}}$) – just before the final decoding step, where the network has refined its scene understanding.
+By constraining intermediate features to be consistent, we encourage the student model to align with the teacher’s representations without relying on explicit pseudo-labels.
+However, excessive weighting of the feature similarity loss may destabilize training, potentially leading to trivial solutions where the student merely replicates the teacher’s features instead of learning meaningful representations.
+3.6 Teacher Fusion
+Unlike BEV segmentation, which includes dynamic classes, all categories in online mapping are stationary.
+This key distinction, combined with the fact that sensor data is collected in sequential frames with known ego vehicle motion, enables novel strategies not commonly explored in other domains.
+We propose to leverage this by fusing pseudo-labels from the teacher across multiple nearby frames.
+As illustrated in Fig.
+1, predictions from both past and future frames can be transformed into the current coordinate frame using the known ego motion.
+By selecting only the most confident teacher predictions from any of these frames, we provide the student with a more complete and accurate map representation.
+This approach is particularly beneficial for improving predictions in regions far from the ego vehicle, where accuracy tends to degrade [13].
+We also explore feature fusion by averaging overlapping BEV feature representations ($\tilde{\mathbf{z}}_{\text{bev}}$) across multiple frames.
+To ensure that additional frames contribute meaningful new information, we randomly sample them at varying distances relative to the current frame.
+This strategy mitigates issues with strictly time-based selection, such as: ego vehicle being stationary, which provides no new information, and the vehicle moving at high speeds, leading to minimal spatial overlap between frames.
+4 Experiments
+Table 1: Performance (mIoU) of various augmentations.
+Using multiple augmentations yields strong results, significantly outperforming no augmentations.
+We also note that CutOut [12] works better than CamDrop [22] for online mapping.
+{NiceTabular}
+[…]
+Photom. CamDrop CutOut BEVDrop $2.5\%$ $10\%$
+$12.1^{\pm 0.6}$ $21.3^{\pm 0.2}$
+$13.3^{\pm 0.6}$ $32.0^{\pm 0.1}$
+✓ ✓ $21.9^{\pm 1.0}$ $31.9^{\pm 0.1}$
+✓ ✓ $24.4^{\pm 1.2}$ $33.9^{\pm 0.2}$
+✓ ✓ ✓ $\textbf{25.5}^{\pm 0.7}$ $\textbf{34.3}^{\pm 0.1}$
+Figure 2: No thresholding, soft MixMatch [2], and hard FixMatch[41] thresholding utilising $2.5\%$ and $10\%$ of the labels. Baseline, mean and std., use neither thresholding nor sharpening.
+To explore the effectiveness of SSL for online mapping, we describe the experimental setup in Sec.
+4.1 and conduct comprehensive ablation studies in Sec.
+4.2 to evaluate the contribution of each component.
+Building on these results, we further evaluate the performance in Sec.
+4.3.
+4.1 Experimental Setup
 We employ a ResNet50 [20] as image encoder $f^{\text{enc}}_{\theta}$, LSS [39] for lifting features to BEV $f^{\text{lift}}_{\theta}$, and a lightweight convolutional bottleneck decoder inspired by [19] as $f^{\text{dec}}_{\theta}$.
@@
 To account for the stochasticity in the results from individual runs, we either report the mean and standard deviation of the mIoU for three runs with the same hyperparameters or multiple nearby values of the hyperparameter.
-Empirically, we note that our setup is robust to a range of unlabelled weights $\omega_{\text{cls}}$ spanning $0.25$ to $5.0$.
-We choose to follow [14] and use $\omega_{\text{cls}}=1$.
-We further follow [21] and linearly ramp-up $\omega_{\text{cls}}$, and when applicable $\omega_{\text{feat}}$, over the first third of the training.
+We choose to follow [15] and use unlabelled weight $\omega_{\text{cls}}=1$.
+We follow [22] and linearly ramp-up $\omega_{\text{cls}}$, and when applicable $\omega_{\text{feat}}$, over the first third of the training.
 Unless stated otherwise, no thresholding or sharpening is used in the experiments.
+For the CutOut augmentation, we remove $25\%$ of the image, and for BEV feature dropout (BEVDrop) we follow [22, 52] and drop $50\%$ of the features.
 The models, unless stated otherwise, are trained and evaluated using the Argoverse 2 (AV2) dataset [49] with the geographically disjoint splits proposed by [28] using surround-view camera images as input and use the provided HD maps as labels.
 We generally experiment using two utilisation rates of labelled data, namely $2.5\%$ and $10\%$, corresponding to $18$ and $70$ labelled sequences respectively.
-IV-B Augmentations
-TABLE I: Performance (mIoU) of various augmentations.
-Augmentation Label utilisation
-Photom. CamDrop CutOut BEVDrop $2.5\%$ $10\%$
-$12.1^{\pm 0.6}$ $21.3^{\pm 0.2}$
-$13.3^{\pm 0.6}$ $32.0^{\pm 0.1}$
-$21.9^{\pm 1.0}$ $31.9^{\pm 0.1}$
-$24.4^{\pm 1.2}$ $33.9^{\pm 0.2}$
+4.2 Exploration
+Table 2: mIoU for sharpening methods using threshold $0.6$. Neither soft MixMatch [2] nor hard thresholding FixMatch[41] show significantly impact over using no thresholding
+{NiceTabular}
+[…]
+Thr. Hard $2.5\%$ $10\%$
 $25.5^{\pm 0.7}$ $34.3^{\pm 0.1}$
-Empirically we observe that using a high probability of applying the augmentation yields better performance and choose always to apply the strong augmentations to the student.
-For the CutOut augmentation, we find that removing $25\%$ of the image yields good results and is thus used for all experiments.
-For BEV feature dropout (BEVDrop) we follow [21, 48] and drop $50\%$ of the features.
-The results in Tab.
-I demonstrate that increasing the difficulty of the task for the student model is crucial also for online mapping.
+$25.1^{\pm 1.3}$ $33.9^{\pm 0.1}$
+✓ ✓ $25.4^{\pm 0.8}$ $34.6^{\pm 0.1}$
+We explore different SSL approaches outlined in Sec. 3 to show their effectiveness for segmentation-based online mapping.
+Augmentations: We evaluate various strong augmentation strategies for the student model, as discussed in Sec.
+3.3.
+The results in Sec.
+4 demonstrate that increasing the difficulty of the task for the student model is crucial also for online mapping.
 With $10\%$ label utilisation, Photometric augmentation alone is sufficient for significant model improvement.
@@
 CutOut, yields better performance than dropping an entire camera and masking its field of view in the loss, as done with CamDrop.
-IV-C Sharpening and Thresholding
-Fig.
-2 illustrates the effects of sharpening and thresholding.
-As shown on the left, supervising with only the most confident pseudo-labels does not yield any additional performance boost.
-Using a low threshold can marginally improve performance, and we opt to use $0.6$ as the threshold in further experiments.
-Tab.
-II displays the results over multiple runs using hard and soft labels.
-We note that there is no significant change in performance using sharpening, which differs from other domains where sharpening has shown to be very effective, usually with high thresholds [37, 48, 44].
-While applying sharpening with a temperature parameter can improve performance, right column in Fig.
-2, the gains are marginal and introduce an additional hyperparameter.
-Figure 2: Thresholding and sharpening utilising $2.5\%$ and $10\%$ of the labels. Baseline refers to using only strong augmentations.
-TABLE II: mIoU for sharpening methods using threshold $0.6$.
-Technique Label utilisation
-Thr. Hard $2.5\%$ $10\%$
-$25.5^{\pm 0.7}$ $34.3^{\pm 0.1}$
-$25.1^{\pm 1.3}$ $33.9^{\pm 0.1}$
-$25.4^{\pm 0.8}$ $34.6^{\pm 0.1}$
-IV-D Feature Consistency
-We compare the performance using Mean Squared Error (MSE) and Cosine Similarity (COS) losses with $\omega_{\text{cls}}=1$ and varying feature loss weight $\omega_{\text{feat}}$.
+Sharpening and Thresholding: The effects of sharpening and thresholding are illustrated in Fig.
+2.
+As shown to the left, supervising with the most confident pseudo-labels does not yield significant performance improvements.
+However, using a low confidence threshold provides a slight boost in performance, and we set the threshold to 0.6 for subsequent experiments.
+The right plot shows the impact of sharpening.
+While sharpening can lead to minor improvements, the gains are marginal and introduce an additional hyperparameter, making it less appealing for our setting.
+Sec.
+4.2 presents results across multiple runs, comparing hard and soft label supervision.
+Overall, these techniques have limited impact on online mapping, contrasting findings in other domains where sharpening has been highly effective—typically when used alongside high thresholds [41].
+Feature Consistency: We compare the performance using Mean Squared Error (MSE) and Cosine Similarity (COS) losses with $\omega_{\text{cls}}=1$ and varying feature loss weight $\omega_{\text{feat}}$.
 In Fig.
-3 the results are displayed using features from before (Pre) or after (Post) the BEV decoder $f^{\text{dec}}_{\theta}$ as inputs to $\ell_{\text{feat}}$.
+3 the results are displayed using features from before or after the BEV decoder $f^{\text{dec}}_{\theta}$ as inputs to $\ell_{\text{feat}}$.
 When utilizing $2.5\%$ of the labels the results are noisy, and no clear benefit from the additional supervision is observed across the setups.
-However, for the $10\%$ label utilisation, COS consistency provides a slight performance gain and demonstrates greater stability than the MSE loss.
-While prior BEV segmentation works opted to use MSE [55, 14], we believe COS to be the better choice.
-Figure 3: Feature Similarity using Mean Square Error and Cosine Similarity with varying weights.
-Baseline refers to using only strong augmentations.
-Pre and Post refers to using $\mathbf{z}_{\text{bev}}$ or $\mathbf{\tilde{z}}_{\text{bev}}$ respectively.
-IV-E Teacher Fusion
-To enhance pseudo-labels, we fuse teacher predictions from multiple samples.
-We use two additional samples, as we empirically found that adding more samples does not provide further benefits.
+However, for the $10\%$ label utilisation, late COS consistency provides a slight performance gain and demonstrates greater stability than the MSE loss.
+While prior BEV segmentation works opted to use MSE [60, 15], we believe COS to be the better choice for our setting.
+Figure 3: Feature Similarity using Mean Square Error following [60] and Cosine Similarity following [46] varying $\omega_{\text{feat}}$.
+Baseline refers to no feature similarity while early and late refers to using BEV features before or after BEV processing $f^{\text{dec}}_{\theta}$ in Fig.
+1.
+Teacher Fusion: To enhance pseudo-labels, we propose to fuse teacher predictions from multiple samples.
+We use two additional samples, as we empirically found, see Sup.
+mat.
+Sec.
+6, that adding more samples does not provide further benefits but makes training slower due to the additional inference costs.
 Therefore, we focus on examining the influence of the maximum distance to these additional samples.
@@
 While the simple probability fusion strategy yields strong results, exploring more advanced fusion techniques is a promising avenue for future research.
-Figure 4: Teacher Fusion across varying maximum ranges from the current sample. Baseline refers using only strong augmentations.
-IV-F Combining Components
-Building on the above results, we further explore how these components interact in SSL for online mapping.
+Figure 4: Teacher Fusion across varying maximum ranges from the current sample. Fusing the probabilities yields consistent improvements over the Baseline without any fusion.
+Combining Components: Building on the above results, we further explore how these components interact in SSL for online mapping.
 Starting with the teacher-student setup as a Core we incrementally add components and evaluate their impact on performance.
@@
 The results in Tab.
-III demonstrate that combining components leads to further increased performance.
+3 demonstrate that combining components leads to further increased performance.
 Adding strong augmentations to the student yields a significant performance boost, as does further incorporating teacher fusion.
@@
 These results underscore the importance of conducting ablation studies across different labelled data ratios, as the impact of various modules can vary depending on the amount of labelled data.
-TABLE III: mIoU when incrementally adding SSL components.
-Label Util. Core +Augs +Fusion +Featsim +Thr. +Hard
+Table 3: mIoU when incrementally adding SSL components.
+Lbl. Util. Core +Augs +Fusion +Featsim +Thr. +Hard
 $2.5\%$ $12.1^{\pm 0.6}$ $25.5^{\pm 0.7}$ $28.2^{\pm 0.7}$ $29.0^{\pm 0.6}$ $29.4^{\pm 0.3}$ $29.3^{\pm 1.1}$
 $10\%$ $21.3^{\pm 0.2}$ $34.3^{\pm 0.1}$ $36.4^{\pm 0.2}$ $36.4^{\pm 0.5}$ $36.8^{\pm 0.2}$ $36.7^{\pm 0.3}$
-IV-G Varying Label Utilisation
-We are also interested in how our proposed SSL method scales with varying amounts of labelled data.
-Using the parameters from Sec.
-IV-F, our proposed method includes strong augmentations, teacher fusion, Cosine feature similarity, and thresholding.
-Tab.
-IV displays performance when using both unlabelled and labelled data against using only the labelled data.
-Each setup is repeated three times with different data folds.
+4.3 Putting SSL to the Test
+Building on the design choices made above, we aim to further verify the strengths of our proposed SSL scheme.
+Using the parameters from Fig.
+4, our proposed method includes strong augmentations, teacher fusion, cosine feature similarity, and thresholding.
 As before, we report the mean and standard deviation; however, in this case, we select the best-performing model on the validation set and evaluate it on the test set.
-Besides conducting experiments on AV2 [45], we also use NuScenes [3], using the same hyperparameters except for learning rate and weight decay.
-As expected, performance decreases as the amount of labelled data is reduced.
-Here, we experiment using $100\%$ of the data, corresponding to $700$ sequences, down to only $18$ sequences at $2.5\%$ utilisation.
-Using the unlabelled data and our proposed setup yields significant performance gains across all labelled data utilisation ratios.
-In the lowest setting, relying solely on the labelled sequences results in poor performance, with mIoU dropping as low as $7.5$ for AV2 and $5.6$ for nuScenes.
-However, incorporating unlabelled data significantly boosts performance to $28.9$ and $15.5$ mIoU for the two datasets respectively.
-Using only $10\%$ of the labels, the performance gap between using the unlabelled data and not is reduced from $29.6$ mIoU to $3.5$ mIoU on AV2, and from $12$ mIoU to $3.4$ mIoU on NuScenes.
-TABLE IV: Test performance (mIoU) for multiple label utilisations for both Argoverse 2 (AV2) and NuScenes (NuSc).
+The performance is compared by utilizing both unlabelled and labelled data with the SSL-scheme versus using only the labelled data in a supervised setting.
+First, we compare our SSL scheme against swapping the backbone for a vision foundation model.
+Next, we evaluate the robustness of our method across different online mapping architectures.
+We then examine its scalability with varying amounts of labelled data and demonstrate the method’s effectiveness for domain adaptation.
+Comparing with pre-training: We substitute the ImageNet pre-trained ResNet50 with DINOv2 [36] vision foundation model, which leverages pre-training on a substantially larger dataset.
+Although DINOv2’s size makes it impractical for vehicular deployment, it demonstrates performance improvements in Tab.
+4.
+Nevertheless, a significant performance gap persists when comparing DINOv2 without SSL to our SSL approach.
+Notably, integrating our SSL scheme with the DINOv2 backbone eliminates this gap, underscoring the effectiveness of our approach.
+Table 4: mIoU when when using DINOv2 backbone adding SSL.
+Backbone SSL Label utilisation
+$2.5\%$ $10\%$
+ResNet50 $7.5^{\pm 0.9}$ $10.6^{\pm 0.2}$
+DINOv2 $11.3^{\pm 0.6}$ $17.3^{\pm 0.4}$
+ResNet50 $29.4^{\pm 0.3}$ $36.8^{\pm 0.2}$
+DINOv2 $28.7^{\pm 0.9}$ $33.2^{\pm 0.2}$
+Robustness to model architecture: We apply our proposed SSL setup to state-of-the-art online mapping models.
+As shown in Fig.
+5, various architectures—including CVT [59], IPM [19], GKT [8], and LSS [39] – all benefit significantly from the SSL scheme.
+LSS achieves the highest overall performance, with a $3.9$x and $3.5$x boost when $2.5\%$ and $10\%$ of the data have labels, respectively.
+While e.g.
+CVT performs slightly worse overall, it sees even greater relative improvements, with performance increasing by $4.8$x and $4.6$x under the same label utilizations.
+Figure 5: Multiple SOTA online mapping methods benefit from the SSL-scheme compared to the supervised training.
+Scalability:
+(a) Argoverse 2
+(b) nuScenes
+Figure 6: Test performance (mIoU) across utilising various amount of the labelled nuScenes data.
+All data is used as unlabelled for all runs.
+Training with unlabelled data using the SSL-scheme consistently improves performance compared to only using supervised training across all label utilisations.
+Figure 7: Qualitative results with varying amounts of labelled data.
+Predictions are shown for models trained with $10\%$, $25\%$, and $100\%$ of the labelled data, using both standard supervised training and our proposed SSL approach.
+Figure 8: Adapting to new cities for AV2.
+Adding unlabelled sequences with the SSL-scheme from the target domain enhances performance in the target city.
+When adapting to Pittsburgh, the gap is almost completely closed to using all target labels.
+We conduct experiments on AV2 [49] and nuScenes [3], with the same hyperparameters except for learning rate and weight decay.
+Our experiments range from using the labels from the full dataset, corresponding to $700$ sequences, down to just $18$ sequences at $2.5\%$ utilisation.
+Figure 6 shows how our proposed SSL method and its supervised counterpart scale with varying amounts of labelled data.
+For the supervised models, performance decreases rapidly as the amount of (labelled) data is reduced.
+In the most label-scarce setting, having access to $2.5\%$ of the labels, models trained solely on labelled sequences perform poorly, with mIoU dropping to $7.5$ for AV2 and $5.6$ for nuScenes.
+Incorporating unlabelled data with our proposed approach yields substantial performance boosts across all labelled data utilisation ratios.
+Leveraging unlabelled data with the SSL-scheme dramatically improves performance to $28.9$ and $15.5$ mIoU for the respective datasets.
+If $10\%$ of the data has labels, our method achieves a $3.5$x performance boost compared to only using the labelled data.
+This narrows the gap to a fully supervised model that uses all labels, to just $3.5$ mIoU on AV2.
+For nuScenes, the performance boost is $1.5$x, and the remaining gap to when all labels have been used is reduced to $3.5$ mIoU.
+To provide visual context for these quantitative improvements, Fig.
+7 shows predictions from models trained only on labelled data against those trained with both labelled and unlabelled data.
+The predictions are noticeably more stable and reliable when using our semi-supervised method, aligning with the quantitative improvements reported above.
+City Adaptation: We assess our method’s ability to adapt to new cities using the AV2 Far Extrapolation splits from [28] and the Boston-to-Singapore split from [35] for nuScenes.
+Our approach utilises all labelled data from the source cities while gradually incorporating unlabelled data from the target cities, ensuring geographic separation from the test set.
+As shown in Fig.
+8 and Fig.
+9 for AV2 and nuScenes, respectively, we find that even without labelled data from the target city, performance improves significantly by utilizing its unlabelled data.
+The domain gap between Boston and Singapore in nuScenes is substantial, making adaptation more challenging than in AV2, where all data comes from US cities.
+Notably, performance improves as more unlabelled sequences from the target domain are added, suggesting that additional unlabelled data could further enhance results.
+For smaller domain gaps, such as adapting to Pittsburgh in AV2, our method nearly matches the performance of a fully supervised model, with only a $0.5$ mIoU difference.
+For larger domain gaps, such as Boston to Singapore, our method boosts mIoU from $11.1$ to $17.0$ when using $100$ unlabelled sequences from Singapore.
+However, a substantial gap remains in cases of large domain shifts, posing a key challenge for future research.
+Figure 9: City adaptation from Boston to Singapore on nuScenes.
+While performance naturally is lower than using all Singapore labels, incorporating unlabelled sequences using the SSL-scheme helps improve the model’s performance when labelled data is only available from Boston.
+5 Conclusion
+This paper draws inspiration from SSL techniques in other domains, demonstrating their effectiveness for segmentation-based online mapping.
+We show that applying photometric, CutOut, and BEVDrop augmentations in a teacher-student setup already performs well.
+When combined with our proposed multi-step fusion, and feature similarity constraints we show an additional performance boost.
+Using the proposed SSL-scheme, we show that incorporating unlabelled data substantially improves performance on both AV2 and nuScenes compared to relying solely on labelled data.
+For example, if $10\%$ of the data has labels, the SSL-approach to leverage unlabelled data achieves a $3.5$x performance boost compared to only using the labelled data, narrowing the gap to a fully supervised model that uses all labels to just $3.5$ mIoU.
+We also show strong generalization to unseen cities and believe that additional unlabelled data could further enhance the domain adaptation results.
+Still, when adapting to Pittsburgh (AV2), incorporating purely unlabelled target-domain data reduces the performance gap from $5$ to just $0.5$ mIoU.
+In summary, efficiently leveraging both labelled and unlabelled data is crucial for advancing online mapping.
+While this work does not explore all existing SSL methods, it provides a foundation for understanding key SSL methods in this context and highlights their potential.
+We hope this study inspires further research in teacher fusion strategies, self-supervised pre-text tasks, and broader semi-supervised learning approaches to enhance online mapping performance further.
+Acknowledgements
+We want to express our sincere gratitude to Adam Tonderski, Georg Hess, William Ljungbergh, Carl Lindström, Ji Lan, and Erik Stenborg for their continuous support and invaluable discussions.
+This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
+Computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at NSC Berzelius and C3SE Alvis partially funded by the Swedish Research Council through grant agreement no.
+2022-06725.
+Supplementary Material
+6 Exact numbers
+For completeness, we also show the exact performance numbers for the results shown in Figures in Sec.
+4.
+For thresholds and temperatures, the numbers are displayed in Tab.
+5 and Tab.
+6 for $2.5\%$ of the labels, respectively.
+For $10\%$, we refer to Tab.
+7 and Tab.
+8.
+In addition to these, the performance for feature similarity loss with different weights is shown in Tab.
+9 and Tab.
+10 for $2.5\%$ and $10\%$ label utilisation, respectively.
+The validation performance for teacher fusion over various ranges is provided in Tab.
+11 and Tab.
+12 for $2.5\%$ and $10\%$ label utilisation.
+We also experiment with the amount of samples to use in addition to the center sample in Tab.
+13.
+For the test performance across multiple label utilisations, the numbers for both Argoverse 2 (AV2) and nuScenes are shown in Tab.
+14.
+Finally, Tab.
+15 provides the performance for city adaptation for both Argoverse 2 and nuScenes when varying the amount of unlabelled data.
+Table 5: Validation performance (mIoU) for different thresholds on Argoverse 2 $2.5\%$ of the labels utilised.
+Thresholds $0.55$ $0.6$ $0.65$ $0.7$ $0.75$ $0.8$ $0.85$ $0.9$
+Thr. $24.1$ $26.8$ $25.9$ $23.6$ $25.7$ $23.6$ $24.6$ $22.3$
+Thr.+Hard $24.4$ $26.6$ $25.2$ $25.5$ $24.2$ $25.0$ $23.7$ $24.6$
+Table 6: Validation performance (mIoU) for using sharpening with different temperatures on Argoverse 2 $2.5\%$ of the labels utilised.
+Temperature $0.05$ $0.1$ $0.25$ $0.5$ $0.75$ $0.95$
+No Thr. $27.6$ $25.3$ $27.4$ $27.1$ $27.0$ $26.1$
+Thr. $24.3$ $26.2$ $27.4$ $25.3$ $25.7$ $26.2$
+Thr.+Hard $24.2$ $25.1$ $25.3$ $26.3$ $26.9$ $24.4$
+Table 7: Validation performance (mIoU) for different thresholds on Argoverse 2 $10\%$ of the labels utilised.
+Thresholds $0.55$ $0.6$ $0.65$ $0.7$ $0.75$ $0.8$ $0.85$ $0.9$
+Thr. $33.6$ $33.9$ $33.4$ $33.3$ $33.0$ $32.4$ $32.1$ $30.9$
+Thr.+Hard $34.8$ $34.7$ $34.4$ $33.7$ $33.3$ $32.3$ $31.4$ $31.4$
+Table 8: Validation performance (mIoU) for using sharpening with different temperatures on Argoverse 2 $10\%$ of the labels utilised.
+Temperature $0.05$ $0.1$ $0.25$ $0.5$ $0.75$ $0.95$
+No Thr. $35.0$ $34.8$ $34.9$ $34.7$ $33.8$ $34.6$
+Thr. $34.2$ $35.3$ $34.8$ $34.2$ $32.9$ $33.5$
+Thr.+Hard $34.6$ $34.8$ $34.4$ $34.9$ $33.9$ $34.1$
+Table 9: Validation performance (mIoU) for feature similarity loss with different weights on Argoverse 2 $2.5\%$ of the labels utilised.
+Weight $0.05$ $0.1$ $0.25$ $0.5$ $1.0$ $1.5$ $2.0$ $2.5$
+MSE-pre $24.3$ $25.4$ $26.0$ $22.2$ $23.3$ $20.0$ $23.1$ $22.5$
+MSE-post $24.3$ $22.5$ $22.5$ $23.1$ $24.3$ $22.5$ $23.4$ $24.1$
+COS-pre $25.3$ $25.3$ $24.8$ $25.0$ $23.4$ $24.9$ $24.6$ $22.9$
+COS-post $23.4$ $25.1$ $23.4$ $23.6$ $25.5$ $26.1$ $24.4$ $25.4$
+Table 10: Validation performance (mIoU) for feature similarity loss with different weights on Argoverse 2 $10\%$ of the labels utilised.
+Weight $0.05$ $0.1$ $0.25$ $0.5$ $1.0$ $1.5$ $2.0$ $2.5$
+MSE-pre $33.7$ $32.8$ $32.3$ $32.5$ $32.4$ $30.0$ $28.8$ $24.1$
+MSE-post $33.6$ $33.4$ $32.3$ $31.9$ $31.2$ $29.3$ $28.0$ $28.0$
+COS-pre $34.4$ $33.8$ $34.4$ $34.4$ $34.1$ $33.8$ $34.4$ $33.2$
+COS-post $34.0$ $34.8$ $35.4$ $35.1$ $34.4$ $34.8$ $33.4$ $31.4$
+Table 11: Validation performance (mIoU) for teacher fusion over various ranges on Argoverse 2 $2.5\%$ of the labels utilised.
+Ranges $10$ $15$ $20$ $25$ $30$ $35$ $40$
+Probs. $27.7$ $27.7$ $27.2$ $27.6$ $28.4$ $28.7$ $26.8$
+Probs.+Thr. $25.5$ $25.9$ $27.3$ $28.2$ $26.9$ $27.5$ $26.7$
+Feats. $24.1$ $25.6$ $24.8$ $23.6$ $23.1$ $25.8$ $26.2$
+Feats.+Thr. $26.5$ $26.6$ $27.0$ $26.5$ $26.1$ $26.7$ $25.6$
+Table 12: Validation performance (mIoU) for teacher fusion over various ranges on Argoverse 2 $10\%$ of the labels utilised.
+Ranges $10$ $15$ $20$ $25$ $30$ $35$ $40$
+Probs. $36.2$ $35.5$ $36.5$ $36.1$ $35.5$ $35.8$ $35.9$
+Probs.+Thr. $35.9$ $36.2$ $36.4$ $35.4$ $35.7$ $35.8$ $36.0$
+Feats. $34.0$ $34.5$ $34.2$ $34.4$ $34.3$ $34.0$ $34.4$
+Feats.+Thr. $34.2$ $34.0$ $34.8$ $34.1$ $34.4$ $35.4$ $34.6$
+Table 13: Multi-step Fusion with varying numbers of additional samples within a 20m range using $10\%$ of labels. The best results are achieved with 4 additional samples.
+Additional samples
+$2$ $4$ $6$
+Probs $34.6$ $35.0$ $30.2$
+Feats $32.4$ $32.8$ $26.1$
+Table 14: Test performance (mIoU) for multiple label utilisations for both Argoverse 2 (AV2) and nuScenes (nuSc).
 +Unlbld Label utilisation
@@
 $\mathbf{15.5}^{\pm 1.2}$ $\mathbf{22.0}^{\pm 1.6}$ $\mathbf{25.1}^{\pm 1.9}$ $\mathbf{27.4}^{\pm 1.8}$ $27.8^{\pm 2.8}$-
-IV-H City Adaptation
-We are interested in determining whether the proposed method can be used to adapt the model to a new city.
-To investigate this, we evaluate the model’s performance using the AV2 Far Extrapolation splits proposed by [27] and the Boston to Singapore split used in [33] for NuScenes.
-For all runs, we use all labelled data from the source cities while gradually increasing the amount of unlabelled data from the target cities, ensuring that this data remains geographically disjoint from the test set.
-As reported in Tab.
-V, we find that, despite having no labelled data from the target city, we can significantly improve performance in that city by leveraging unlabelled data from it.
-In the AV2 dataset, our approach nearly matches the performance of a model trained using all target labels not included in the test set.
-E.g., when adapting to Pittsburgh, the performance difference between using labelled data and our method is only $0.5$ mIoU.
-However, it is noteworthy that the domain gap between the cities in AV2 is smaller since all cities are located in the United States.
-For NuScenes, the domain gap between Boston and Singapore is substantial, making it a more challenging problem.
-Yet our proposed method increases the mIoU from $11.1$ to $17.0$ using $100$ unlabelled sequences from Singapore.
-TABLE V: Performance (mIoU) for city adaptaion for both Argoverse 2 and NuScenes when increasing amount of unlabelled data.
+Table 15: Performance (mIoU) for city adaptaion for both Argoverse 2 and nuScenes when increasing amount of unlabelled data.
 Folds # unlabelled target sequences Using target
@@
 nuSc Boston $\rightarrow$ Singapore $11.1$ $12.6$ $15.3$ $17.0$ $17.7$ $27.0$
-V Conclusion
-This paper draws inspiration from semi-supervised learning (SSL) techniques in other domains, demonstrating how to apply them to online mapping.
-For instance, we have shown that applying Photometric, CutOut, and BEV feature dropout augmentations to the student network significantly improves performance in a teacher-student setup.
-Further, we show that our proposed multi-step fusion for improving the teachers’ pseudo-labels yields an additional performance boost and that the Cosine similarity loss can help improve learning.
-Combining the methods explored in this paper, performance can be significantly increased when also utilising unlabelled data for both Argoverse 2 and NuScenes compared to only using the labelled data.
-For instance, merely using $10\%$ of the labels, the performance gap between using the unlabelled data and not is reduced from $29.6$ mIoU to $3.5$ mIoU on Argoverse 2, and from $12$ mIoU to $3.4$ mIoU on NuScenes.
-We have also shown that our method can be used for city adaptation, where we have labels for one city but not for the other.
-For example, when adapting from Boston to Singapore, performance is increased by $6.6$ mIoU when unlabelled data from Singapore is included during training.
-In summary, efficient use of both labelled and unlabelled data is crucial for advancing the field of online mapping.
-While we could not explore all existing SSL methods, this work takes a step toward understanding the strengths and weaknesses of core SSL methods in this context, aiming to inspire further research in this area.
-Future research could explore more rigorous teacher fusion over multiple samples or incorporate self-supervision tasks for further advancement.
-Acknowledgements We want to express our sincere gratitude to Adam Tonderski, Georg Hess, William Ljungbergh, Carl Lindström, and Ji Lan for their continuous support and invaluable discussions.
-This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
-Computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at NSC Berzelius and C3SE Alvis partially funded by the Swedish Research Council through grant agreement no.
-2022-06725.